# Non-Relativistic Limits of Bosonic and Heterotic Double Field Theory

### v1

@@ Abstract
The known stringy non-relativistic (NR) limit of the universal NS-NS sector of supergravity has a finite Lagrangian due to non-trivial cancellations of divergent parts coming from the metric and the $B$-field. We demonstrate that in Double Field Theory (DFT) and generalised geometry these cancellations already happen at the level of the generalised metric, which is convergent in the limit $c\rightarrow\infty$. On the level of bosonic DFT this implies that the NR limit can be taken before imposing the strong constraint and breaking the duality symmetry. At the (finite) leading order this reproduces the Non-Riemannian formulation of DFT. Assuming a convergent generalised metric for heterotic DFT, we derive a novel $c$-expansion for heterotic supergravity which is, by construction, O$(D,D)$-invariant.
@@ 1 Introduction
The effect of T-duality as a fundamental symmetry of NRST has been studied in several works. This symmetry maps a space-like longitudinal direction to a null one, underlying a discrete lightcone quantization in the NRST. In this work, we rewrite the NR supergravity in the language of Double Field Theory (DFT) [15, 16, 17, 18, 19, 20] and generalised geometry, which are (double) geometries in which the low energy limit of ST can be written in a T-duality invariant way (see [21, 22, 23] and the second lecture of [24] for reviews on this topic). For the bosonic case, the generalised metric ^$\hat{\cal H}(\hat{g},\hat{B})$ contains divergences in its components, but when one considers the $B$-field expansion (1.2), the components of ^$\hat{\cal H}(\hat{g},\hat{B})$ are all convergent, i.e.,
@@
which in turn can be understood as a non-Riemannian generalised metric, like in the original proposal of [25] (see also [26, 27, 28, 29, 30]). This can be summarized in the following way:
supergravity ($\begin{array}[]{ccc}\textrm{supergravity }(\hat{g},\hat{B},\hat{\Phi})&\longleftrightarrow&\textrm{bosonic DFT }(\hat{\cal H},\hat{d})\\ \downarrow&&\downarrow\\ \textrm{NR limit }(\tau,e,b,\phi)&\longleftrightarrow&\textrm{O}(D,D)\textrm{-invariant NR limit }({\cal H}^{(0)},d)\end{array}$ (1.9)
In this diagram, the downwards arrow denotes the limit $c\rightarrow\infty$ while the double arrows mean
$\rightarrow$ : to double the coordinates and construct T-duality invariant multiplets,
$\leftarrow$ : to break the duality group.
A similar observation was also made in the case of exceptional field theory for Newton-Cartan limits of $M$-theory [31].
@@
We start embedding the proposal given by Bergshoeff and Romano [13] in DFT. In this case the generalised metric has divergent contributions
${\cal H}=c^{4}{\cal H}^{(4)}+c^{2}{\cal H}^{(2)}+{\cal H}^{(0)}+\mathcal{O}\left(\frac{1}{c^{2}}\right)\,.$ (1.10)
@@
In section 2 we give a quick review to the supergravity formulation of NRST. We set up our notation and we present the fields, the symmetry transformations and the action. Then in section 3 we reformulate the bosonic supergravity in the standard language of DFT. We start by constructing the generalised frame, and from it we construct the generalised metric. After considering the $B$-field expansion, we arrive to convergent expansion for the generalised metric. We compare these results with the non-Riemannian formulation of DFT in the last part of this section. Then, in section 4 we follow a similar procedure but including the heterotic gauge field in a duality covariant way. Following the prescription of [13] leads to a finite supergravity but the components of the generalised metric are divergent. Therefore, in the final part of this section we give an alternative proposal for the expansion of heterotic supergravity, which leads to a finite generalised metric. Finally, we conclude with a discussion in section 5.
@@ 3 Bosonic Double Field Theory
Let us split $\hat{A}=(A,A^{\prime})$ where $A,B,...=0,\dots,3$ is a double transversal index and $A^{\prime},B^{\prime},...=4,\dots,2D-1$ then the generalised frame $\hat{E}_{M}{}^{\hat{A}}$ decomposes as follows:
$\hat{E}_{M}{}^{A}=\frac{1}{\sqrt{2}}\left(\begin{matrix}-c{\tau}_{\mu\underline{a}}-\frac{1}{c}\hat{B}_{\rho\mu}{\tau}^{\rho}{}_{\underline{a}}&\frac{1}{c}{\tau}^{\mu}{}_{\underline{a}}\\ c\tau_{\mu\overline{a}}-\frac{1}{c}\hat{B}_{\rho\mu}{}\tau^{\rho}{}_{\overline{a}}&\frac{1}{c}\tau^{\mu}{}_{\overline{a}}\end{matrix}\right)\,,$ (3.1)
and
$\hat{E}_{M}{}^{A^{\prime}}=\frac{1}{\sqrt{2}}\left(\begin{matrix}-{e}_{\mu\underline{a}^{\prime}}-\hat{B}_{\rho\mu}{e}^{\rho}{}_{\underline{a}^{\prime}}&{e}^{\mu}{}_{\underline{a}^{\prime}}\\ e_{\mu\overline{a}^{\prime}}-\hat{B}_{\rho\mu}{}e^{\rho}{}_{\overline{a^{\prime}}}&e^{\mu}{}_{\overline{a}^{\prime}}\end{matrix}\right)\,,$ (3.2)
@@ 3.2 Generalised fluxes and action
These $c$-expansion components satisfy
@@
From (3.34) we learn that both ^$\hat{\cal H}_{MP}$ and ^$\hat{\cal H}^{(0)}_{MP}$ are O$(D,D)$-elements. Moreover, as also the generalised dilaton is finite, $\hat{d}=d$, the DFT action is finite:
@@
and therefore the Lagrangian is well defined as NR limit of DFT.
@@ 3.3 Comparison with non-Riemannian DFT
So far, we have shown that the NR limit of bosonic DFT is consistent, and leads to finite degrees of freedom, transformations and action. A similar analysis has been obtained in the context of Exceptional Field Theory for the membrane Newton-Cartan parameterisation [31]. In the next section, we will study expansions of heterotic supergravity for which such an analysis in terms of heterotic DFT is missing in the literature, so far.
@@ 4.2 An alternative $c$-expansion
    where $F^{(a)}$ is the usual non-Abelian field strength for the sub-leading connection $a$. The gauge transformations take a natural form:
    $\delta_{\lambda}a_{\mu}^{k}$ $=$ $\partial_{\mu}\lambda^{k}+f^{k}{}_{ij}\lambda^{i}a_{\mu}^{j},$ (4.33)
@@ 5 Discussion
The present article continues a series of works [26, 27, 28, 29, 31, 35] that highlight the usefulness of DFT and generalised geometry in understanding NR limits of supergravity. In particular, arranging degrees of freedom of supergravity in NR limits in terms of the generalised metric of DFT makes is tractable how cancellations of divergences happens. The generalised metric also appears naturally in the Hamiltonian formulation of the string worldsheet [36, 37, 38, 15, 16, 39, 40]. This was extended to the heterotic world-sheet in [41, 42] and to general worldvolumes in string and $M$-theory in the context of Exceptional Field Theory in [43, 44, 45, 46, 47]. Hence, one could expect that $c$-expansions in DFT or Exceptional Field Theory are useful when taking the (string or $p$-brane) NR-limits of these worldvolume theories. As for the point particle in Newton-Cartan geometry or the string in String Newton-Cartan geometry, it is expected that a more general expansion, including $\frac{1}{c^{2}}$-corrections of the $\tau$, is necessary for this [5].
@@ Acknowledgements
The work of E.L. is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland. The work of D.O. is part of project No. 2022/45/P/ST2/03995 co-funded by the National Science Centre (NCN) and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 945339.

### v3

@@ Abstract
The known stringy non-relativistic (NR) limit of the universal NS-NS sector of supergravity has a finite Lagrangian due to non-trivial cancellations of divergent parts coming from the metric and the $B$-field. We demonstrate that in Double Field Theory (DFT) and generalised geometry these cancellations already happen at the level of the generalised metric, which is convergent in the limit $c\rightarrow\infty$, implying that the NR limit can be imposed before solving the strong constraint. We present the $c$-expansion of the generalised metric, which reproduces the Non-Riemannian formulation of DFT at the (finite) leading order, and the $c$-expansion of the generalised frame, which contains divergences. We also extend this approach to the non-Abelian gauge field of Heterotic DFT assuming a convergent expansion for the O$(D,D+n)$ generalised metric. From this proposal, we derive a novel $c$-expansion for the bosonic part of the heterotic supergravity which is, by construction, compatible with O$(D,D)$-symmetry.
@@ 1 Introduction
The effect of T-duality as a fundamental symmetry of NRST has been studied in several works. This symmetry maps a space-like longitudinal direction to a null one, underlying a discrete lightcone quantisation in the NRST. In this work, we rewrite the NR supergravity in the language of Double Field Theory (DFT) [15, 16, 17, 18, 19, 20] and generalised geometry, which are (double) geometries in which the low energy limit of ST can be written in a T-duality invariant way (see [21, 22, 23] and the second lecture of [24] for reviews on this topic). For the bosonic case, the generalised metric ^$\hat{\cal H}(\hat{g},\hat{B})$ contains divergences in its components, but when one considers the $B$-field expansion (1.2), the components of ^$\hat{\cal H}(\hat{g},\hat{B})$ are all convergent, i.e.,
@@
which is a so-called non-Riemannian generalised metric [25]. For the flat space, the Gomis-Ooguri background [1], this had been demonstrated in [26] and for arbitrary backgrounds in [27]. The proporties of this Non-Riemmanian limit of DFT have been studied in [28, 29, 27, 30, 26]. Motivated by this result, a similar observation was also made in the case of exceptional field theory for Newton-Cartan limits of M-theory [31].
For the bosonic case, we extend the existing discussion by presenting the $c$-expansion of DFT both in the generalised metric and generalised frame formalism:
supergravity bosonic DFT ($\begin{array}[]{ccccc}\textrm{supergravity }&\longleftrightarrow&\textrm{bosonic DFT }&&\\ (\hat{g},\hat{B},\hat{\Phi})&&(\hat{\cal H},\hat{d})\text{ or }(\hat{E},\hat{d})&&\\ \downarrow&&\downarrow&\searrow&\\ \text{{non}-finite $c$-expansion}&\longleftrightarrow&\text{finite $c$-expansion}&&\text{{non}-finite $c$-expansion}\\ (\hat{g},\hat{B},\hat{\Phi})&&(\hat{\mathcal{H}},\hat{d})&&(\hat{E},\hat{d})\\ \downarrow&&\downarrow&&\downarrow\\ \textrm{finite NR-supergravity action}&\longleftrightarrow&\text{Non-Riemannian DFT}&&\text{finite NR-DFT action}\\ (\tau,e,b,\phi)&&({\cal H}^{(0)},d)&&\end{array}$ (1.14)
In this diagram, the upper downwards arrow denotes the $c$-expansion of the fundamental variables of the theory, wheres the lower downwards-arrow denotes the limit $c\rightarrow\infty$. The $\leftrightarrow$-arrow means
$\rightarrow$ : doubling the coordinates and construct T-duality invariant multiplets,
$\leftarrow$ : breaking the duality group.
The endpoints of this diagram (the first and last row) have been presented before. We argue that the $c$-expansion (the middle row) is also relevant for two reasons:
  • It shows that the $c$-expansion for the generalised frame variables $\hat{E}$ is not finite. This becomes relevant when considering the NR-limit of DFT with $\alpha^{\prime}$-corrections where there is no generalised metric formulation [32] in general. See section 3.4 for more details.
  • $\frac{1}{c^{2}}$-corrections are relevant for the study of the string world-sheet in DFT and generalised geometry. For example, in analogy to the point particle one also has to consider $\frac{1}{c^{2}}$-corrections to $\tau$, the string Bargmann field [5, 33, 34].
@@
We start embedding the proposal given by Bergshoeff and Romano [13] in DFT. In this case, the generalised metric has divergent contributions
^$\hat{\cal H}=c^{4}{\cal H}^{(4)}+c^{2}{\cal H}^{(2)}+{\cal H}^{(0)}+\mathcal{O}\left(\frac{1}{c^{2}}\right)\,.$ (1.15)
@@
In section 2 we give a quick review to the supergravity formulation of NRST. We set up our notation and we present the fields, the symmetry transformations and the action. Then in section 3 we reformulate the bosonic supergravity in the standard language of DFT. We start by constructing the generalised frame, and from it we construct the generalised metric. After considering the $B$-field expansion, we arrive to convergent expansion for the generalised metric. Then we compare these results with the non-Riemannian formulation of DFT, finding agreement. At the end of this section we discuss the implications of the divergent terms in the generalized frame formulation for the higher-derivative structure of DFT. Then, in section 4 we follow a similar procedure but including the heterotic gauge field in a duality covariant way. Following the prescription of [13] leads to a finite supergravity but the components of the generalised metric are divergent. Therefore, in the final part of this section we give an alternative proposal for the expansion of heterotic supergravity, which leads to a finite generalised metric. Finally, we conclude with a discussion in section 5.
@@ 3 Bosonic Double Field Theory
Firstly, we compute the generalised frame corresponding to the above expansion of $\hat{e}$ and $\hat{B}$. This shows that the generalised frame does not possess a finite $c$-expansion, similar to the frame $\hat{e}$ in the supergravity expansion. For this $\hat{A}=(A,A^{\prime})$ with $A,B,...=0,\dots,3$ is a double transversal index and $A^{\prime},B^{\prime},...=4,\dots,2D-1$ then the generalised frame $\hat{E}_{M}{}^{\hat{A}}$ decomposes as follows:
$\hat{E}_{M}{}^{A}$ $=c{E^{(1),A}_{M}}+\frac{1}{c}{E^{(-1),A}_{M}}\,,$ (3.1)
$\text{with}\quad{E^{(1),A}_{M}}$ $=\frac{1}{\sqrt{2}}\left(\begin{matrix}-\tau_{\mu\underline{a}}-{\tau_{\mu}}^{c}\epsilon_{bc}\delta^{b}_{\underline{a}}&0\\ \tau_{\mu\overline{a}}-{\tau_{\mu}}^{c}\epsilon_{bc}\delta^{b}_{\overline{a}}&0\end{matrix}\right),\quad{E^{(-1),A}_{M}}=\frac{1}{\sqrt{2}}\left(\begin{matrix}-b_{\rho\mu}{\tau^{\rho}}_{\underline{a}}&{\tau^{\mu}}_{\underline{a}}\\ -b_{\rho\mu}{\tau^{\rho}}_{\overline{a}}&{\tau^{\mu}}_{\overline{a}}\end{matrix}\right)$ (3.2)
and the finite
$\hat{E}_{M}{}^{A^{\prime}}=\frac{1}{\sqrt{2}}\left(\begin{matrix}-{e}_{\mu\underline{a}^{\prime}}-\hat{B}_{\rho\mu}{e}^{\rho}{}_{\underline{a}^{\prime}}&{e}^{\mu}{}_{\underline{a}^{\prime}}\\ e_{\mu\overline{a}^{\prime}}-\hat{B}_{\rho\mu}{}e^{\rho}{}_{\overline{a^{\prime}}}&e^{\mu}{}_{\overline{a}^{\prime}}\end{matrix}\right)=\frac{1}{\sqrt{2}}\left(\begin{matrix}-{e}_{\mu\underline{a}^{\prime}}-b_{\rho\mu}{e}^{\rho}{}_{\underline{a}^{\prime}}&{e}^{\mu}{}_{\underline{a}^{\prime}}\\ e_{\mu\overline{a}^{\prime}}-b_{\rho\mu}{}e^{\rho}{}_{\overline{a^{\prime}}}&e^{\mu}{}_{\overline{a}^{\prime}}\end{matrix}\right)\,.$ (3.3)
@@ 3.2 Generalised fluxes and action
At leading order, this had been demonstrated in [26] for flat space (Gomis-Ooguri background [1]) for arbitrary backgrounds in [27]. The proporties of this Non-Riemmanian limit of DFT have been studied in [28, 29, 27, 30, 26].
The $c$-expansion components satisfy
@@
From (3.35) we learn that both ^$\hat{\cal H}_{MP}$ and ^$\hat{\cal H}^{(0)}_{MP}$ are O$(D,D)$-elements. Moreover, as also the generalised dilaton is finite, $\hat{d}=d$. Hence, the DFT action is finite [26, 27]
@@
and therefore the Lagrangian is well defined as NR limit of DFT. A similar analysis has been obtained in the context of Exceptional Field Theory for the membrane Newton-Cartan parameterisation [31]. In section 4, we will study expansions of heterotic supergravity for which such an analysis in terms of heterotic DFT is missing in the literature, so far.
@@ 3.3 Comparison with non-Riemannian DFT
With this, we have shown that the complete $c$-expansion of bosonic DFT in the generalised metric formalism takes very similar form the Non-Riemmanian formulation. Only $H^{\mu\nu}$ is modified and with that the property that $x$ lies in the kernel of $H$.
3.4 Comments about higher-derivative terms
One important application of DFT is the construction of the four derivative terms of the bosonic and heterotic supergravity starting from an O$(D,D)$ action which schematically we can write as
$S=\int d^{2D}X\ \left({\cal R}({\hat{F}})+a{\cal R}^{-}(\hat{F})+b{\cal R}^{+}(\hat{F})\right),$ (3.44)
and the explicit form of the action in terms of the generalized fluxes can be found in [35]. In order to find $\alpha^{\prime}$-corrections after parametrising the fundamental fields we need to impose
$(a,b)=\begin{cases}(-1,-1)&\mbox{bosonic string \, , }\\ (-1,0)&\mbox{heterotic string \, ,}\\ (0,0)&\mbox{type II strings \, .}\end{cases}$ (3.45)
As initially discussed in [32], there exists a tension between the generalised metric and frame formalism beyond the two-derivative level. Particularly, only the combination $(a,b)=(-1,1)$ can be written in terms of the generalised metric and the generalised dilaton. This theory does not correspond to a string theory formulation and it was originally constructed in [36] and further explored in [37, 38]. This implies that invoking the results of this section, only the HSZ theory enjoys, in principle, a finite Lagrangian using the parameterisation (3.29)-(3.31). This means that for other choices of the parameters apart from the HSZ theory the divergences in the generalised frame (and also in the generalised fluxes) we cannot ensure, in principle, a finite structure for ${\cal R}^{-}$ and ${\cal R}^{+}$. In order to construct higher-derivative corrections from the double geometry for the NR bosonic and heterotic supergravity, we then expect modifications in the generalised Green-Schwartz transformations and, consequently, in the four derivative Lagrangian.
In the next section, we return to the two-derivative case and focus on the heterotic formulation of DFT [39].
@@ 4.2 An alternative $c$-expansion
    where $\nabla^{(a)}_{\mu}=\partial_{\mu}+[\cdot,a_{\mu}]$ is the covariant derivative and $F^{(a)}$ the usual non-Abelian field strength for the sub-leading connection $a$. The gauge transformations take a natural form,
    $\delta_{\lambda}a_{\mu}^{k}$ $=$ $\partial_{\mu}\lambda^{k}+f^{k}{}_{ij}\lambda^{i}a_{\mu}^{j}=\nabla^{(a)}_{\mu}\lambda^{k},$ (4.33)
@@ 5 Discussion
The present article continues a series of works [28, 29, 27, 30, 31, 42] that highlight the usefulness of DFT and generalised geometry in understanding NR limits of supergravity. In particular, arranging degrees of freedom of supergravity in NR limits in terms of the generalised metric of DFT makes is tractable how cancellations of divergences happens. The generalised metric also appears naturally in the Hamiltonian formulation of the string worldsheet [43, 44, 45, 15, 16, 46, 47]. This was extended to the heterotic world-sheet in [48, 49, 50] and to general worldvolumes in string and $M$-theory in the context of Exceptional Field Theory in [51, 52, 53, 54, 55]. Hence, one could expect that $c$-expansions in DFT or Exceptional Field Theory are useful when taking the (string or $p$-brane) NR-limits of these worldvolume theories. As for the point particle in Newton-Cartan geometry or the string in String Newton-Cartan geometry, it is expected that a more general expansion, including $\frac{1}{c^{2}}$-corrections of the $\tau$, is necessary for this [5].
@@ Acknowledgements
We are grateful to Chris D.A. Blair for comments on the first version and helpful remarks on the development of the finiteness of the non-relativistic limit of bosonic DFT in the literature. E.L. thanks Mariana Graña for enlighting discussions in the first stages of this work during his visit Université Paris-Saclay covered by the ERC Consolidator Grant 772408-Stringlandscape. E.L. is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland. The work of D.O. is part of project No. 2022/45/P/ST2/03995 co-funded by the National Science Centre (NCN) and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 945339.